OpenStudy (espex):

Use of an integrating factor: "Find the general solution of the given differential equation. State an interval on which the general solution is defined. \[cos(x) \frac{dy}{dx} + ysin(x) = 1\]

OpenStudy (anonymous):

Is this one separable?

OpenStudy (espex):

I started by dividing out the cos(x) to get \[\frac{dy}{dx} + ytan(x) = \frac{1}{cos(x)}\] Then created an integrating factor: \[IF = e^{\int\limits_{}^{}P(x) dx} \rightarrow e^{\int\limits_{}^{}\tan(x) dx}\]

OpenStudy (espex):

I do not believe it is, though I did not try and separate it.

OpenStudy (espex):

After integrating I had \[e^{-\ln|\cos(x)|}\] which I multiplied both sides of the equation by. \[e^{-\ln|\cos(x)|} \frac{dy}{dx} + e^{-\ln|\cos(x)|} ytan(x) = e^{-\ln|\cos(x)|} \frac{1}{\cos(x)}\]

OpenStudy (espex):

Not exactly sure where to go from here.

OpenStudy (anonymous):

Refer to the attachment, a screen capture .

OpenStudy (turingtest):

@eSpeX you still there?

OpenStudy (turingtest):

\[e^{-\ln\cos x}=\frac1{\cos x}=\sec x\]

OpenStudy (espex):

@TuringTest Thank you, I was pulling my hair out because of that silly sec.